# Towards sub-milliarcsecond astrometric precision using seeing-limited imaging

### v1

@@ 10.1 Bootstrap
To estimate the precision and consistency of our results, we perform a Bootstrap test (Efron 1982). We split the dataset into two subsets (i.e., even and odd indices), run the pipeline separately for each subset, and compare the results. The Bootstrap test is conducted for multiple scenarios, each incorporating different components of the astrometric model.
Table 2 provides the Bootstrap rms for different algorithm configurations. Figure 9 compares the proper motion measurements between the two subsets for Algorithm 6 (see Table 2), and Figure 10 shows the 2D proper motion differences as a function of magnitude. The comparison is shown for two groups of sources: (1) sources with a single-epoch precision of $<10$ mas (orange) and (2) sources with a single-epoch precision between $10$ and $40$ mas (blue). Additionally, the comparison indicates a precision level of approximately [^7] Reduced to $0.07$ mas year-1 if one considers that the number of observations were reduced by a factor of two in the Bootstrap test. $0.1$ mas year-1. The best theoretical precision expected for this dataset is about $0.04$ mas ($\approx 10/\sqrt{3200}$), where 10 mas represents the per-epoch precision, and 3200 is the assumed number of epochs. Consequently, the precision inferred from the bootstrap analysis is roughly a factor of 2–3 worse than the theoretical expectation.
Figure 9: Bootstrap test for proper motion in the BLG17K0103 field. Figure 9(a) (Figure 9(b)) shows the difference in proper motion along the x-axis (y-axis) between the two datasets, plotted as a function of the proper motion from dataset 1. Figure 9(c) (9(d)) shows the histogram of the x-axis (y-axis) proper motion difference between the two data sets. The red colour denotes sources with single-epoch precision between 10 and 40 mas, while the violet colour indicates sources with single-epoch precision better than 10 mas.
@@ 10.2 Comparison with Gaia DR3
Due to the high density of stars in the Galactic bulge, Gaia measurements in this region are sparse and even missing for many stars. Nevertheless, we attempted to compare our results with Gaia. We cross-matched our sources with Gaia DR3 (Prusti et al., 2016; Gaia Collaboration, 2022), excluding outliers (28 out of 105 sources; see §6.1) as well as sources with ruwe$>$1.3 (5 additional sources, of which two were already among the outliers). Figure 11 shows the differences between the Gaia and our measured proper motion as a function of the stellar right ascension and declination. Although there is a clear linear relation between the Gaia and KMTNet-based measurements, as expected, these relations are not one-to-one, and they show a clear offset. The reason for this is that while Gaia is measuring proper motion in a global coordinate system (ICRS), while the KMTNet measurements are done with respect to a relative coordinate system. Furthermore, while the mean stellar environment (and proper motion) may change along the KMTNet fields, this may introduce some complicated relations between the Gaia and KMTNet proper motion systems. To take this into account, we fitted the difference in proper motion to
@@
The values of the parameters are field-dependent. Figure 12 compares the Gaia and KMTNet proper motions after this fit was performed. In Table 2, we also provide the rms compared to Gaia proper motion measurements after fitting this transformation. We note that the median of the Gaia proper motion errors (accuracy) for the sources we use for the comparison is 0.073 mas year-1 (0.12 mas year-1) for declination (right ascension), which are comparable with our proper motion precision, as can be seen in Figure 13 which shows the ratio between the difference of proper motion (Gaia-KMT) and the Gaia proper motion accuracy. The observed error ratio, approximately $\mathord{\sim}3$, may also be indicative of an underestimation of the Gaia uncertainties in the Galactic Bulge. This overestimation of the GAIA errors in the bulge is also supported by Vasiliev & Baumgardt (2021) and Luna et al. (2023).
@@
Figure 12: Comparison of proper motions in the 17K0103 field with Gaia DR3. Figures 12(a) and 12(b) show proper motion differences in right ascension and declination (KMTNet vs. Gaia DR3), respectively, as a function of the Gaia DR3 proper motion. Figures 12(c) and 12(d) display histograms of proper motion differences in right ascension and declination after outlier removal. The legend indicates the standard deviation of proper motion differences under specific conditions. The bin size is 0.1 mas yr-1.
@@
The same selection criteria and quality cuts described earlier for BLG17K0103 were applied, including limiting the sample to sources brighter than $I\mathord{<}18.5$, with nearest-neighbour distances greater than 5 pixels, and excluding observations with an airmass $>1.5$ or FWHM $>4.5$ pixels (1.8″).
@@
Figure 14: Comparison of proper motions, between CTIO and SAAO observatories, in the BLG15M0306 field. Figure 14(a) (14(b)) shows the x-axis (y-axis) proper motion difference between the two telescopes, CTIO-SAAO, as a function of the proper motion measured in CTIO. Figure 14(c) (14(d)) shows the histogram of the x-axis (y-axis) proper motion differences between the two observatories. The blue colour represents sources with single-epoch precision better than 10 mas, while the red colour represents sources with precision between 10 and 40 mas.
@@ 11 Conclusion
We presented a pipeline for precise astrometric measurements of ground-based observations, aiming to achieve milliarcsecond-level astrometric precision necessary for detecting isolated compact objects via astrometric microlensing. The pipeline is available online888 Will be added with dependencies on https://github.com/EranOfek/AstroPack.
The pipeline was demonstrated on approximately $\mathord{\sim}6500$ images from the KMTNet BLG17K0103 field, successfully measuring and correcting for systematic effects such as Differential Chromatic Refraction (DCR) and the Annual effect (see §8.1 and §8.2). To evaluate the precision of the proper motions derived, we use the Bootstrap technique and compare the results with Gaia DR3 data. The Bootstrap test indicates a precision of $\mathord{\sim}0.1$ mas/year, while the comparison with Gaia DR3 yields proper motion precisions of $\mathord{\sim}0.3$ mas year-1. The precision as estimated compared to Gaia DR3, is a factor of $\mathord{\sim}3$ higher than the precision estimated from the Bootstrap test. This may suggest that either the precision of the Gaia-DR3 proper motion measurements in the bulge is underestimated (which is supported by Vasiliev & Baumgardt 2021; Luna et al. 2023), and/or there is an additional systematic error in our measurements.
@@ Appendix A PSF - Multivariate t-distribution
×$\quad\times\left[1+\nu^{-1}\left(\textbf{x}-\mu\right)^{T}\Sigma^{-1}\left(\textbf{x}-\mu\right)\right]^{-(\nu+p)/2}$ (38)
@@
Data Availability
The raw KMTNet images used in this work are not publicly available but can be accessed upon request through the Korea Astronomy and Space Science Institute (KASI) and subject to their data-sharing policies. Derived data products and astrometric catalogues generated by our pipeline will be shared on reasonable request to the corresponding author. The astrometry pipeline code will be accessible via GitHub.

### v2

@@ 10.1 Bootstrap
To estimate the precision and consistency of our results, we perform a Bootstrap test (efron1982jackknife). We split the dataset into two subsets (i.e., even and odd indices), run the pipeline separately for each subset, and compare the results. The Bootstrap test is conducted for multiple scenarios, each incorporating different components of the astrometric model.
Table 2 provides the Bootstrap rms for different algorithm configurations. Figure 10 compares the proper motion measurements between the two subsets for Algorithm 6 (see Table 2), and Figure 11 shows the 2D proper motion differences as a function of magnitude. The comparison is shown for two groups of sources: (1) sources with a single-epoch precision of <10<10 mas (orange) and (2) sources with a single-epoch precision between $10$ and $40$ mas (blue). Additionally, the comparison indicates a precision level of approximately [^7] Reduced to $0.07$ mas year-1 if one considers that the number of observations were reduced by a factor of two in the Bootstrap test. $0.1$ mas year-1. The best theoretical precision expected for this dataset is about $0.04$ mas ($\approx 10/\sqrt{3200}$), where 10 mas represents the per-epoch precision, and 3200 is the assumed number of epochs. Consequently, the precision inferred from the bootstrap analysis is roughly a factor of 2–3 worse than the theoretical expectation.
Figure 10: Bootstrap test for proper motion in the BLG17K0103 field. Figure 10(a) (Figure 10(b)) shows the difference in proper motion along the x-axis (y-axis) between the two datasets, plotted as a function of the proper motion from dataset 1. Figure 10(c) (10(d)) shows the histogram of the x-axis (y-axis) proper motion difference between the two data sets. The violet colour denotes sources with single-epoch precision between 10 and 40 mas, while the red colour indicates sources with single-epoch precision better than 10 mas.
@@ 10.2 Comparison with Gaia DR3
Due to the high density of stars in the Galactic bulge, Gaia measurements in this region are sparse and even missing for many stars. Nevertheless, we attempted to compare our results with Gaia. We cross-matched our sources with Gaia DR3 (GAIA_mission_prusti2016gaia; GAIA+2022yCat_GAIA_DR3_MainSourcesCatalog), excluding outliers (28 out of 105 sources; see §6.1) as well as sources with ruwe>>1.3 (5 additional sources, of which two were already among the outliers). Figure 12 shows the differences between the Gaia and our measured proper motion as a function of the stellar right ascension and declination. Although there is a clear linear relation between the Gaia and KMTNet-based measurements, as expected, these relations are not one-to-one, and they show a clear offset. The reason for this is that while Gaia is measuring proper motion in a global coordinate system (ICRS), while the KMTNet measurements are done with respect to a relative coordinate system. Furthermore, while the mean stellar environment (and proper motion) may change along the KMTNet fields, this may introduce some complicated relations between the Gaia and KMTNet proper motion systems. To take this into account, we fitted the difference in proper motion to
@@
The values of the parameters are field-dependent. Figure 13 compares the Gaia and KMTNet proper motions after this fit was performed. In Table 2, we also provide the rms compared to Gaia proper motion measurements after fitting this transformation. We note that the median of the Gaia proper motion errors (accuracy) for the sources we use for the comparison is 0.073 mas year-1 (0.12 mas year-1) for declination (right ascension), which are comparable with our proper motion precision, as can be seen in Figure 14 which shows the ratio between the difference of proper motion (Gaia-KMT) and the Gaia proper motion accuracy. The observed error ratio, approximately $\mathord{\sim}3$, may also be indicative of an underestimation of the Gaia uncertainties in the Galactic Bulge. This underestimation of the GAIA errors in the bulge is also supported by vasiliev2021gaia_error_underestimation and luna2023Gaia_astrometry_underestimation.
@@
Figure 13: Comparison of proper motions in the 17K0103 field with Gaia DR3. Figures 13(a) and 13(b) show proper motion differences in right ascension and declination (KMTNet vs. Gaia DR3), respectively, as a function of the Gaia DR3 proper motion. Figures 13(c) and 13(d) display histograms of proper motion differences in right ascension and declination after outlier removal. The legend indicates the standard deviation of proper motion differences under specific conditions. The bin size is 0.1 mas yr-1.
@@
10.2.1 Comparison of Absolute Positions with Gaia
In addition to the proper-motion comparison described above, we also examined the consistency of our measured absolute positions with those from Gaia DR3. The comparison was performed in the same manner as the proper-motion analysis, using the affine transformations defined in Equations 35 and 36.
A direct comparison between the absolute positions derived from our relative astrometric solution and those from Gaia DR3, shown in Figure 15, reveals that the positional scatter (rms $\approx 4$ mas) is roughly twice as large as expected from the formal uncertainties ($\approx 1.5$ mas, estimated by multiplying the proper-motion rms by the 7 yr time baseline). Such an excess is naturally expected for a relative astrometric solution, since the global reference frame is free to drift with respect to the inertial system. The excess scatter likely originates from residual systematics inherent to the relative nature of our solution. Because the astrometry is solved relative to the mean reference frame rather than the ICRS, slow variations in the effective zero point of the frame, driven by seasonal atmospheric conditions or colour-dependent refraction, can shift the local coordinate system with respect to Gaia. In particular, the annual effect (§8.2), which reflects residual colour- and season-dependent refraction terms not fully captured by the basic chromatic model, may introduce correlated positional offsets that vary on yearly timescales. Such offsets, combined with small residual geometric distortions or long-term drift of the fiducial frame, can naturally lead to the observed excess scatter.
To test whether colour-dependent effects contribute to the observed scatter, we divided the sources into two colour bins, ${\rm Bp-Rp}<3.5$ and ${\rm Bp-Rp}\geq 3.5$, and repeated the position comparison for each subset (Figure 15). The resulting distributions show comparable dispersions, with only a modest difference between the blue and green subsamples, suggesting that while chromatic systematics may be present, they are not the dominant source of the observed positional scatter.
It is important to note that even systematics that are partially absorbed by the affine or chromatic terms can propagate into the reference positions $(x_{0},y_{0})$, since these parameters are derived from the same global fit. Consequently, any residual frame-dependent offsets, whether seasonal, chromatic, or geometric, manifest as excess scatter in the absolute position comparison with Gaia, while having a much smaller impact on the relative proper motions. Therefore, the proper-motion comparison with Gaia remains reliable, although it should be interpreted as relative rather than tied to an absolute reference frame.
Figure 15: Comparison of absolute positions in the 17K0103 field with Gaia DR3. The upper and lower panels show histograms of positional differences in right ascension and declination (KMTNet – Gaia), respectively. Sources are divided by colour, with blue and green histograms representing stars with ${\rm Bp-Rp}<3.5$ and ${\rm Bp-Rp}\geq 3.5$, respectively. The legend indicates the standard deviation of the positional differences for each subsample. Only sources with rms<10<10 mas are included. The bin size is 1 mas.
@@ 10.2.1 Comparison of Absolute Positions with Gaia
The same selection criteria and quality cuts described earlier for BLG17K0103 were applied, including limiting the sample to sources brighter than $I\mathord{<}18.5$, with nearest-neighbour distances greater than 5 pixels, and excluding observations with an airmass >1.5>1.5 or FWHM >4.5>4.5 pixels (1.8″).
@@
Figure 16: Comparison of proper motions, between CTIO and SAAO observatories, in the BLG15M0306 field. Figure 16(a) (16(b)) shows the x-axis (y-axis) proper motion difference between the two telescopes, CTIO-SAAO, as a function of the proper motion measured in CTIO. Figure 16(c) (16(d)) shows the histogram of the x-axis (y-axis) proper motion differences between the two observatories. The blue colour represents sources with single-epoch precision better than 10 mas, while the red colour represents sources with precision between 10 and 40 mas.
@@ 11 Conclusion
We presented a pipeline for precise astrometric measurements of ground-based observations, aiming to achieve milliarcsecond-level astrometric precision necessary for detecting isolated compact objects via astrometric microlensing. The pipeline is available online [^8] Code availability: The Lizoo pipeline and installation instructions are available at https://github.com/noamse/Lizoo.
The pipeline was demonstrated on approximately $\mathord{\sim}6500$ images from the KMTNet BLG17K0103 field, successfully measuring and correcting for systematic effects such as Differential Chromatic Refraction (DCR) and the Annual effect (see §8.1 and §8.2). At the bright end (mag $\lesssim$17), we reach a per-epoch astrometric precision of $\mathord{\sim}$5 mas. In addition, we quantified short-cadence astrometric precision by binning the residual time series over 5–20 day cadences. For bright sources, this stacking reduces the weighted rms to $\sim$2 mas per coordinate (Figure 7), close to the precision required for astrometric microlensing, and the consistent improvement across cadences indicates that the residual scatter scales down with temporal stacking.
To evaluate the precision of the proper motions derived, we use the Bootstrap technique and compare the results with Gaia DR3 data. The Bootstrap test indicates a precision of $\mathord{\sim}0.1$ mas/year, while the comparison with Gaia DR3 yields proper motion precisions of $\mathord{\sim}0.3$ mas year-1. The precision as estimated compared to Gaia DR3, is a factor of $\mathord{\sim}3$ higher than the precision estimated from the Bootstrap test. This may suggest that either the precision of the Gaia-DR3 proper motion measurements in the bulge is underestimated (which is supported by vasiliev2021gaia_error_underestimation; luna2023Gaia_astrometry_underestimation), and/or there is an additional systematic error in our measurements.
@@ Acknowledgements
Data Availability
The raw KMTNet images used in this work are not publicly available but can be accessed upon request through the Korea Astronomy and Space Science Institute (KASI) and subject to their data-sharing policies. Derived data products and astrometric catalogues generated by our pipeline will be shared on reasonable request to the corresponding author. The Lizoo astrometric pipeline code is available on GitHub at https://github.com/noamse/Lizoo.
@@ Appendix A PSF - Multivariate t-distribution
$\quad\times\left[1+\nu^{-1}\left(\textbf{x}-\mu\right)^{T}\Sigma^{-1}\left(\textbf{x}-\mu\right)\right]^{-(\nu+p)/2}$ (38)